# Flattening Hierarchies with Policy Bootstrapping

### v1

@@ 4.4 Hierarchies perform test-time policy bootstrapping
HIQL separately trains a low-level subpolicy on goals sampled from states at most $k$ steps into the future, reaping all the benefits of policy training with nearby goals. Similar to other goal-conditioned hierarchical methods [28, 24], it then uses the high-level policy to predict optimal subgoals between the current state and the goal at test time, and “bootstraps” by using the subgoal-conditioned action distribution as an estimate for the full goal-conditioned policy.
@@ 5 Subgoal Advantage-Weighted Policy Bootstrapping
We now seek to unify the above insights into an objective to learn a single, flat goal-reaching policy without the additional complexity of HRL. Following the bootstrapping perspective, a direct analogue to hierarchies would use the subpolicy to construct training targets, regressing the full goal-conditioned policy towards a target subpolicy conditioned on the output of a subgoal generator $P\left(a\mid s,\pi^{h}(w\mid s,g)\right)$. This approach, taken by RIS [Figure 1], still inherits the full complexity of hierarchical policies and then some: it requires learning a subgoal generator, a subpolicy, and an additional flat policy.
@@ 5.1 Hierarchical RL as inference
To eliminate this additional machinery, we adopt the view of GCRL as probabilistic inference [22]. In this framing, the bilevel objectives for HIQL’s hierarchical policy and the KL bootstrapping term for RIS’s flat policy can be derived from the same inference problem with different choices of variational posterior. Our main insight is that the expectation over generated subgoals can be expressed as an expectation over the dataset distribution with an advantage-based importance weight, yielding our SAW objective. We present an abridged version below and leave the full derivation to Appendix C.
Similar to previous work [1], we cast the infinite-horizon, discounted GCRL formulation as an inference problem by constructing a probabilistic model via the likelihood function $p\left(U=1\mid\tau,\{w\},g\right)\propto\exp\left(\beta\sum_{t=0}^{\infty}\gamma^{t}A\left(s_{t},w_{t},g\right)\right)$, where the binary variable $U$ can be interpreted as the event of reaching the goal $g$ as quickly as possible from state $s_{t}$ by passing through subgoal state $w$. The subgoal advantage is defined as $A(s_{t},w,g)=-V(s_{t},g)+\gamma^{k}V(w,g)+\sum_{t^{\prime}=t}^{k-1}r(s_{t^{\prime}},g)$, where $w\sim p^{\mathcal{D}}(s_{t+k}\mid s_{t})$. In practice, we follow HIQL and simplify the advantage estimate to $V(w,g)-V(s_{t},g)$, i.e., the progress towards the goal achieved by reaching $w$.
Without loss of generality (since we can represent any flat Markovian policy simply by setting $\pi^{h}(\cdot\mid s,g)$ to a point distribution on $g$), we use an inductive bias on the subgoal structure of GCRL to consider prior distributions $\pi$ of a factored hierarchical form
@@
The distinctions between hierarchical approaches like HIQL and non-hierarchical approaches such as RIS and SAW begin with our choice of variational posterior. For the former, we would consider similarly factored distributions, whereas for the latter, we use a flat policy $\pi_{\theta}(\tau\mid g)$ that factors as
$\pi_{\theta}(\tau\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\pi_{\theta}(a_{t}\mid s_{t},g),$
assuming that the dataset policies are Markovian. We also introduce a variational posterior $q(\{w\}\mid g)$ which factors over a sequence of waypoints $\{w\}=\{w_{0},w_{1},\ldots\}$ as
$q(\{w\}\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\pi^{\mathrm{sub}}(a_{t}\mid s_{t},w_{t})q(w_{t}\mid s_{t},g).$
@@
[…]
$\geq\mathbb{E}_{\pi_{\theta}(\tau\mid g),q(\{w\}\mid g),p(g)}\log\left[\frac{p(U=1,\tau,\{w\}\mid g)}{\pi_{\theta}(\tau\mid g)q(\{w\}\mid g)}\right]=\mathcal{J}(\pi_{\theta},q).$
Expanding distributions according to their factorizations, dropping terms that are independent of the variationals, and rewriting the discounted sum over time as an expectation over the (unnormalized) discounted stationary state distribution $\mu_{\pi}(s)=\sum_{t=0}^{\infty}\gamma^{t}p(s_{t}=s\mid\pi)$ results in the final objective
$\mathcal{J}(\theta,q)=\mathbb{E}_{\mu(s),p(g)}\big{[}\mathbb{E}_{q(w\mid s,g)}\left[A(s,w,g)\right]$ $-\mathbb{E}_{\pi_{\theta}(a\mid s,g)}\left[D_{\mathrm{KL}}(q(w\mid s,g)\|\pi^{h}(w\mid s,g))\right]$
$-\mathbb{E}_{q(w\mid s,g)}\left[D_{\mathrm{KL}}(\pi_{\theta}(a\mid s,g)\|\pi^{\ell}(a\mid s,w))\right]\big{]},$ (5)
where we optimize an approximation of $\mathcal{J}$ by sampling from the dataset distribution $\mu_{\mathcal{D}}(s)$ [37, 1, 35].
@@
7:     Train target subpolicy: $\psi\leftarrow\psi-\lambda\nabla_{\psi}\mathcal{J}_{\mathrm{AWR}}(\psi)$ with $(s_{t},a,w)\sim p^{\mathcal{D}}$ [Equation 3]
@@ 5.2 Eliminating the subgoal generator
Both RIS and HIQL directly parameterize $q(w\mid s,g)$ with a subgoal generator and optimize according the first line of Equation 5.1, which can be solved analytically and results in an objective similar to AWR [Equation 3] when projected into the space of parameterized policies
$\mathcal{J}(q)$ $=\mathbb{E}_{\mu(s),p(g)}\left[\mathbb{E}_{q(w\mid s,g)}[A(s,w,g)]-\mathbb{E}_{\pi_{\theta}(a\mid s,g)}\left[D_{\mathrm{KL}}(q(w\mid s,g)\|\pi^{h}(w\mid s,g))\right]\right]$
$\propto\mathbb{E}_{\mu(s),p(g)}\left[\exp(A(s,w,g))\log q(w\mid s,g)\right].$
Then, RIS trains a flat policy $\pi_{\theta}$ using the third term in Equation 5.1, which is an expectation over the subgoals generated by $q$. Our key insight is that, rather than learning a generative model to approximate $p(w\mid s,g,U=1)$, we can instead use a simple application of Bayes’ rule:
@@
We separately learn an approximation to $\pi^{\ell}(a\mid s,w)$, which we do in practice by training a target subpolicy with AWR in a similar fashion to HIQL (whereas RIS uses a exponential moving average of its online goal-conditioned policy as a target). While we omit this for clarity, a subpolicy AWR term can be incorporated into our objective by introducing another optimality variable $p\left(O_{t}=1\mid\tau,\{w\}\right)$ and a posterior for $\pi^{\ell}$, similar to our derivation for HIQL in Appendix C.
While approximating $p(w\mid s,U=1)$ with $q$ directly and using our importance weight on the dataset distribution are mathematically equivalent, the latter does introduce sampling-based limitations, which we discuss in Appendix A. However, we show empirically that the benefits from lifting the burden of learning a distribution over a high-dimensional subgoal space far outweigh these drawbacks, especially in large state spaces with high intrinsic dimensionality.
@@ 5.3 The SAW objective
The importance weight in Equation 6 allows the policy to bootstrap from subgoals sampled directly from dataset trajectories by ensuring that only subpolicies conditioned on high-advantage subgoals influence the direction of the goal-conditioned policy. We combine our bootstrapping term with an additional learning signal from a (one-step) policy extraction objective utilizing the value function, which improves performance in stitching-heavy environments [Appendix H]. Here, we use one-step AWR [Equation 3], yielding the full SAW objective:
$\mathcal{J}(\theta)=\mathbb{E}_{p^{\mathcal{D}}(s,a,w),p(g)}{\left[e^{\alpha A(s,a,g)}\log\pi_{\theta}(a\mid s,g)-e^{\beta A(s,w,g)}D_{\mathrm{KL}}\left(\pi_{\theta}(\cdot\mid s,g)\|\pi^{\mathrm{sub}}_{\psi}(\cdot\mid s,w)\right)\right]}$ (7)
@@ 6.1 Experimental setup
We select several environments and their corresponding datasets from the recently released OGBench suite [31], a comprehensive benchmark specifically designed for offline GCRL. OGBench provides multiple state-goal pairs for evaluation and datasets tailored to evaluate desirable properties of offline GCRL algorithms, such as the ability to reason over long horizons and stitch across multiple trajectories or combinatorial goal sequences. We use the baselines from the original OGBench paper, which include both one-step and hierarchical state-of-the-art offline GCRL methods. We briefly describe each category of tasks below and baseline algorithms in Appendix D, and encourage readers to refer to the Park et al. [31] for further details.
@@
[…]
Environment Dataset GCBC GCIVL GCIQL QRL CRL HIQL SAW
pointmaze pointmaze-medium-navigate-v0 $9$$\pm 6$ $63$$\pm 6$ $53$$\pm 8$ $82$$\pm 5$ $29$$\pm 7$ $79$$\pm 5$ $\mathbf{97}$$\pm 2$
pointmaze-large-navigate-v0 $29$$\pm 6$ $45$$\pm 5$ $34$$\pm 3$ $\mathbf{86}$$\pm 9$ $39$$\pm 7$ $58$$\pm 5$ $\mathbf{85}$$\pm 10$
pointmaze-giant-navigate-v0 $1$$\pm 2$ 00 ±$\pm 0$ 00 ±$\pm 0$ $\mathbf{68}$$\pm 7$ $27$$\pm 10$ $46$$\pm 9$ $\mathbf{68}$$\pm 8$
antmaze antmaze-medium-navigate-v0 $29$$\pm 4$ $72$$\pm 8$ $71$$\pm 4$ $88$$\pm 3$ $\mathbf{95}$$\pm 1$ $\mathbf{96}$$\pm 1$ $\mathbf{97}$$\pm 1$
antmaze-large-navigate-v0 $24$$\pm 2$ $16$$\pm 5$ $34$$\pm 4$ $75$$\pm 6$ $83$$\pm 4$ $\mathbf{91}$$\pm 2$ $\mathbf{90}$$\pm 3$
antmaze-giant-navigate-v0 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ $14$$\pm 3$ $16$$\pm 3$ ${65}$$\pm 5$ $\mathbf{73}$$\pm 4$
humanoidmaze humanoidmaze-medium-navigate-v0 $8$$\pm 2$ $24$$\pm 2$ $27$$\pm 2$ $21$$\pm 8$ $60$$\pm 4$ $\mathbf{89}$$\pm 2$ $\mathbf{88}$$\pm 3$
humanoidmaze-large-navigate-v0 $1$$\pm 0$ $2$$\pm 1$ $2$$\pm 1$ $5$$\pm 1$ $24$$\pm 4$ $\mathbf{49}$$\pm 4$ ${46}$$\pm 4$
humanoidmaze-giant-navigate-v0 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ $1$$\pm 0$ $3$$\pm 2$ ${12}$$\pm 4$ $\mathbf{35}$$\pm 4$
cube cube-single-play-v0 $6$$\pm 2$ $53$$\pm 4$ ${68}$$\pm 6$ $5$$\pm 1$ $19$$\pm 2$ $15$$\pm 3$ $\mathbf{72}^{*}$$\pm 5$
cube-double-play-v0 $1$$\pm 1$ $36$$\pm 3$ $\mathbf{40}$$\pm 5$ $1$$\pm 0$ $10$$\pm 2$ $6$$\pm 2$ $\mathbf{40}$$\pm 7$
cube-triple-play-v0 $1$$\pm 1$ $1$$\pm 0$ $3$$\pm 1$ 00 ±$\pm 0$ $\mathbf{4}$$\pm 1$ $3$$\pm 1$ $\mathbf{4}$$\pm 2$
scene scene-play-v0 $5$$\pm 1$ $42$$\pm 4$ ${51}$$\pm 4$ $5$$\pm 1$ $19$$\pm 2$ $38$$\pm 3$ $\mathbf{63}$$\pm 6$
visual-antmaze visual-antmaze-medium-navigate-v0 $11$$\pm 2$ $22$$\pm 2$ $11$$\pm 1$ 00 ±$\pm 0$ $\mathbf{94}$$\pm 1$ $\mathbf{93}$$\pm 4$ $\mathbf{95}$$\pm 0$
visual-antmaze-large-navigate-v0 $4$$\pm 0$ $5$$\pm 1$ $4$$\pm 1$ 00 ±$\pm 0$ $\mathbf{84}$$\pm 1$ $53$$\pm 9$ $\mathbf{82}$$\pm 4$
visual-antmaze-giant-navigate-v0 00 ±$\pm 0$ $1$$\pm 1$ 00 ±$\pm 0$ 00 ±$\pm 0$ $\mathbf{47}$$\pm 2$ $6$$\pm 4$ $10$$\pm 2$
visual-cube visual-cube-single-play-v0 $5$$\pm 1$ $60$$\pm 5$ $30$$\pm 5$ $41$$\pm 15$ $31$$\pm 15$ $\mathbf{89}$$\pm 0$ $\mathbf{88}$$\pm 3$
visual-cube-double-play-v0 $1$$\pm 1$ $10$$\pm 2$ $1$$\pm 1$ $5$$\pm 0$ $2$$\pm 1$ $\mathbf{39}$$\pm 2$ $\mathbf{40}$$\pm 3$
visual-cube-triple-play-v0 $15$$\pm 2$ $14$$\pm 2$ $15$$\pm 1$ $16$$\pm 1$ $17$$\pm 2$ $\mathbf{21}$$\pm 0$ $\mathbf{20}$$\pm 1$
visual-scene visual-scene-play-v0 $12$$\pm 2$ $25$$\pm 3$ $12$$\pm 2$ $10$$\pm 1$ $11$$\pm 2$ $\mathbf{49}$$\pm 4$ $\mathbf{47}$$\pm 6$

### v3

@@ 4.4 Hierarchies perform test-time policy bootstrapping
HIQL separately trains a low-level subpolicy on goals sampled from states at most $k$ steps into the future, reaping all the benefits of policy training with nearby goals. Similar to other goal-conditioned hierarchical methods [nachum_data-efficient_2018, levy_learning_2018], it then uses the high-level policy to predict optimal subgoals between the current state and the goal at test time, and “bootstraps” by using the subgoal-conditioned action distribution as an estimate for the full goal-conditioned policy.
@@ 5 Subgoal Advantage-Weighted Policy Bootstrapping
We now seek to unify the above insights into an objective to learn a single, flat goal-reaching policy without the additional complexity of HRL. Following the bootstrapping perspective, a direct analogue to hierarchies would use the subpolicy to construct training targets, regressing the full goal-conditioned policy towards a target action distribution $\pi^{\mathrm{sub}}\left(a\mid s,\pi^{h}(w\mid s,g)\right)$ given by a subpolicy $\pi^{\mathrm{sub}}$ conditioned on optimal subgoals from a subgoal generator, similar to Equation 5.
Deriving such an objective in a principled manner leads us to a key contribution of this work: by viewing hierarchical policy optimization as probabilistic inference over the optimal distribution of subgoals, we can (1) derive SAW, our training-time bootstrapping objective that does not require generative subgoal models; and (2) recover RIS, a related bootstrapping objective that still requires a high-level policy [chane-sane_goal-conditioned_2021]; as well as (3) the bilevel HIQL objective [Equation 4.1]. We present an abridged version below and leave the full derivations to Appendix B.
@@ 5.1 Hierarchical RL as inference
The connection between RL and probabilistic inference is well-established [dayan_using_1996, rawlik_stochastic_2008, levine_variational_2013, levine_reinforcement_2018, abdolmaleki_maximum_2018, peng_advantage-weighted_2019] and begins by constructing a probabilistic model over states, actions, and a binary optimality variable $O_{t}$. The joint likelihood of optimality ($O_{t}=1$ for $t=\{t,\ldots,T\}$) over a trajectory $\tau$ is chosen as $p(O=1\mid\tau)\propto\exp\left(\sum_{t=0}^{\infty}\gamma^{t}r_{t}/\alpha\right)$, where $\alpha$ is a temperature parameter. Intuitively, $O_{t}$ represents the event of maximizing return by choosing an action, i.e., maximizing progress towards the goal in the goal-conditioned case. peng_advantage-weighted_2019 derive a similar form that instead maximizes the expected policy improvement $J(\pi)-J(\pi_{\beta})$ of $\pi$ over a behavior policy $\pi_{\beta}$ in terms of the advantage $A^{\pi_{\beta}}(s,a)=Q^{\pi_{\beta}}(s,a)-V^{\pi_{\beta}}(s)$, which we use both for compatibility with HIQL and for practical improvements in stability. We provide a more thorough introduction to these ideas in Appendix A.
To paint hierarchical policy optimization as an inference problem, we define a variable $U$ over subgoals, where $p\left(U=1\mid\tau,\{w\},g\right)\propto\exp\left(\beta\sum_{t=0}^{\infty}\gamma^{t}A\left(s_{t},w_{t},g\right)\right)$. The binary variable $U$ can be interpreted as the event of reaching the goal $g$ as quickly as possible from state $s_{t}$ by passing through subgoal state $w$. The subgoal advantage is defined as $A(s_{t},w,g)=-V(s_{t},g)+\gamma^{k}V(w,g)+\sum_{t^{\prime}=t}^{k-1}r(s_{t^{\prime}},g)$. In practice, we follow HIQL and simplify the advantage estimate to $V(w,g)-V(s_{t},g)$, i.e., the progress towards the goal achieved by reaching $w$ [Appendix F.1].
Without loss of generality (since we can represent any flat Markovian policy simply by setting $\pi^{h}(\cdot\mid s,g)$ to a point distribution on $g$), we use an inductive bias on the subgoal structure of GCRL to consider policies $\pi$ of a factored hierarchical form
@@
The distinctions between hierarchical approaches like HIQL and non-hierarchical approaches such as RIS and SAW begin with our choice of posterior distribution. For the former, we would consider similarly factored distributions, whereas for the latter, we use a flat posterior $q^{f}(\tau\mid g)$ that factors as
$q^{f}(\tau\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})q^{f}(a_{t}\mid s_{t},g),$
assuming that the dataset policies are Markovian. We also introduce a posterior $q^{h}(\{w\}\mid g)$ which factors over a sequence of waypoints $\{w\}=\{w_{0},w_{1},\ldots\}$ as
$q^{h}(\{w\}\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\pi^{\mathrm{sub}}(a_{t}\mid s_{t},w_{t})q^{h}(w_{t}\mid s_{t},g),$
@@
$\log p_{\pi}(U=1)$ $=\log\int p(g)p_{\pi}(\tau,\{w\}\mid g)p(U=1\mid\tau,\{w\},g)d\{w\}\,d\tau\,dg$
$\geq\mathbb{E}_{q^{f}(\tau\mid g),q^{h}(\{w\}\mid g),p(g)}\log\left[\frac{p(U=1,\tau,\{w\}\mid g)}{q^{f}(\tau\mid g)q^{h}(\{w\}\mid g)}\right]=\mathcal{J}(q,\pi).$
Here, $q^{f}$ and $q^{h}$ serve as auxiliary distributions that allow us to optimize their corresponding parametric policies $\pi_{\theta}$ and $\pi^{h}$ via an iterative expectation-maximization (EM) procedure. Expanding distributions according to their factorizations, dropping terms that are independent of the variationals, and rewriting the discounted sum over time as an expectation over the (unnormalized) discounted stationary state distribution $\mu_{\pi}(s)=\sum_{t=0}^{\infty}\gamma^{t}p(s_{t}=s\mid\pi)$ results in the final objective
$\mathcal{J}(q,\pi)=\mathbb{E}_{\mu(s),p(g)}\big[\mathbb{E}_{q^{h}(w\mid s,g)}\left[A(s,w,g)\right]$ $-\mathbb{E}_{q^{f}(a\mid s,g)}\left[D_{\mathrm{KL}}(q^{h}(w\mid s,g)\|\pi^{h}(w\mid s,g))\right]$
$-\mathbb{E}_{q^{h}(w\mid s,g)}\left[D_{\mathrm{KL}}(q^{f}(a\mid s,g)\|\pi^{\ell}(a\mid s,w))\right]\big],$ (6)
where we optimize an approximation of $\mathcal{J}$ by sampling from the dataset distribution $\mu_{\mathcal{D}}(s)$ [schulman_trust_2015, abdolmaleki_maximum_2018, peng_advantage-weighted_2019].
@@
7:  Train target subpolicy: $\omega\leftarrow\omega-\lambda\nabla_{\omega}\mathcal{J}_{\mathrm{AWR}}(\omega)$ with $(s_{t},a,w)\sim p^{\mathcal{D}}$ [Equation 3]
@@ 5.2 Eliminating the subgoal generator
Converting the KL penalty in the first line of Equation 5.1 to a hard constraint $D_{\mathrm{KL}}(\cdot\|\cdot)\leq\epsilon$ allows us to find the closed form of the optimal posterior $q^{h}(w\mid s,g)\propto\pi^{h}_{\psi}(w\mid s,g)\exp(A(s,w,g))$ [abdolmaleki_maximum_2018, peng_advantage-weighted_2019]. Both RIS and HIQL iteratively minimize the KL divergence between the parametric generative model $\pi^{h}_{\psi}(w\mid s,g)$ and this sample-based posterior $q^{h}$. Then, RIS trains a flat policy $\pi_{\theta}$ using the remaining KL divergence term in the second line. This minimizes the divergence between the flat goal-conditioned policy posterior $q^{f}$ and the subgoal-conditioned policy $\pi^{\ell}$—in expectation over the optimal distribution of subgoals $q^{h}$, as approximated by $\pi^{h}$
$\mathcal{J}^{f}_{\mathrm{RIS}}(\theta)=\mathbb{E}_{\mu(s),q^{h}(w),p(g)}\left[D_{\mathrm{KL}}\left(\pi_{\theta}(a\mid s,g)\|\pi^{\mathrm{sub}}(a\mid s,w)\right)\right].$ (7)
This brings us to the key insight underlying the SAW objective: instead of learning a generative model over the potentially high-dimensional space of subgoals to compute the expectation over $q^{h}(w)$, we can directly estimate this expectation from dataset samples using a simple application of Bayes’ rule:
@@
We learn an approximation to $\pi^{\ell}(a\mid s,w)$ by training a separate (sub)goal-conditioned policy with AWR in a similar fashion to HIQL, whereas RIS uses an exponential moving average of the parameters of its full goal-conditioned policy as a target. Since our regression target is a parametric distribution, this conveniently allows us to directly model $q^{f}$ with another parametric policy $\pi_{\theta}(a\mid s,g)$.
While approximating $p(w\mid s,U=1)$ with $q$ directly and using our importance weight on the dataset distribution are mathematically equivalent, the latter does introduce sampling-based limitations, which we discuss in Section 7. However, we show empirically that the benefits from lifting the burden of learning a distribution over a high-dimensional subgoal space far outweigh these drawbacks, especially in large state spaces with high intrinsic dimensionality.
@@ 5.3 The SAW objective
The importance weight in Equation 8 allows the policy to bootstrap from subgoals sampled directly from dataset trajectories by ensuring that only subpolicies conditioned on high-advantage subgoals influence the direction of the goal-conditioned policy [Figure 1]. We combine our bootstrapping term with an additional learning signal from a (one-step) policy extraction objective utilizing the value function, which improves performance in stitching-heavy environments [Appendix I]. Here, we use one-step AWR [Equation 3], yielding the full SAW objective:
$\mathcal{J}(\theta)=\mathbb{E}_{p^{\mathcal{D}}(s,a,w),p(g)}{\left[e^{\alpha A(s,a,g)}\log\pi_{\theta}(a\mid s,g)-e^{\beta A(s,w,g)}D_{\mathrm{KL}}\left(\pi_{\theta}(a\mid s,g)\|\pi^{\mathrm{sub}}(a\mid s,w)\right)\right]}$ (9)
@@ 6.1 Experimental setup
We select several environments and their corresponding datasets from the recently released OGBench suite [park_ogbench_2024], a comprehensive benchmark specifically designed for offline GCRL. OGBench provides multiple state-goal pairs for evaluation and datasets tailored to evaluate desirable properties of offline GCRL algorithms, such as the ability to reason over long horizons and stitch across multiple trajectories or combinatorial goal sequences. We use the baselines from the original OGBench paper, which include both one-step and hierarchical state-of-the-art offline GCRL methods. We briefly describe each category of tasks below and baseline algorithms in Appendix C, and encourage readers to refer to the park_ogbench_2024 for further details.
@@
Table 1: Evaluating SAW on state- and pixel-based offline goal-conditioned RL tasks. We compare our method’s average (binary) success rate ($\%$) against the numbers reported in park_ogbench_2024 across the five test-time goals for each environment, averaged over 88 seeds (44 seeds for pixel-based visual tasks) with standard deviations after the $\pm$ sign. Numbers within $5\%$ of the best value in the row are in bold. Results with an asterisk () use different value learning hyperparameters and are discussed further in Section 6.3.
Environment Dataset GCBC GCIVL GCIQL QRL CRL HIQL RISoff SAW
pointmaze pointmaze-medium-navigate-v0 99 $\pm 6$ $63$ $\pm 6$ $53$ $\pm 8$ $82$ $\pm 5$ $29$ $\pm 7$ $79$ $\pm 5$ $88$ $\pm 6$ $\mathbf{97}$ $\pm 2$
pointmaze-large-navigate-v0 $29$ $\pm 6$ $45$ $\pm 5$ $34$ $\pm 3$ $\mathbf{86}$ $\pm 9$ $39$ $\pm 7$ $58$ $\pm 5$ $63$ $\pm 13$ $\mathbf{85}$ $\pm 10$
pointmaze-giant-navigate-v0 11 $\pm 2$ 0 $\pm 0$ 0 $\pm 0$ $\mathbf{68}$ $\pm 7$ $27$ $\pm 10$ $46$ $\pm 9$ $57$ $\pm 12$ $\mathbf{68}$ $\pm 8$
antmaze antmaze-medium-navigate-v0 $29$ $\pm 4$ $72$ $\pm 8$ $71$ $\pm 4$ $88$ $\pm 3$ $\mathbf{95}$ $\pm 1$ $\mathbf{96}$ $\pm 1$ $\mathbf{96}$ $\pm 1$ $\mathbf{97}$ $\pm 1$
antmaze-large-navigate-v0 $24$ $\pm 2$ $16$ $\pm 5$ $34$ $\pm 4$ $75$ $\pm 6$ $83$ $\pm 4$ $\mathbf{91}$ $\pm 2$ $\mathbf{89}$ $\pm 3$ $\mathbf{90}$ $\pm 3$
antmaze-giant-navigate-v0 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ $14$ $\pm 3$ $16$ $\pm 3$ 65{65} $\pm 5$ 65{65} $\pm 4$ $\mathbf{73}$ $\pm 4$
humanoidmaze humanoidmaze-medium-navigate-v0 88 $\pm 2$ $24$ $\pm 2$ $27$ $\pm 2$ 2121 $\pm 8$ $60$ $\pm 4$ $\mathbf{89}$ $\pm 2$ $73$ $\pm 5$ $\mathbf{88}$ $\pm 3$
humanoidmaze-large-navigate-v0 11 $\pm 0$ 22 $\pm 1$ 22 $\pm 1$ 55 $\pm 1$ $24$ $\pm 4$ $\mathbf{49}$ $\pm 4$ 21{21} $\pm 7$ 46{46} $\pm 4$
humanoidmaze-giant-navigate-v0 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ 11 $\pm 0$ 33 $\pm 2$ 12{12} $\pm 4$ 33 $\pm 2$ $\mathbf{35}$ $\pm 4$
cube cube-single-play-v0 66 $\pm 2$ $53$ $\pm 4$ 68{68} $\pm 6$ 55 $\pm 1$ 1919 $\pm 2$ $44^{*}$ $\pm 9$ $\mathbf{81}^{*}$ $\pm 6$ ${72}^{*}$ $\pm 5$
cube-double-play-v0 11 $\pm 1$ $36$ $\pm 3$ $\mathbf{40}$ $\pm 5$ 11 $\pm 0$ $10$ $\pm 2$ 66 $\pm 2$ $36$ $\pm 4$ $\mathbf{40}$ $\pm 7$
cube-triple-play-v0 11 $\pm 1$ 11 $\pm 0$ 33 $\pm 1$ 0 $\pm 0$ $\mathbf{4}$ $\pm 1$ 33 $\pm 1$ 33 $\pm 2$ $\mathbf{4}$ $\pm 2$
scene scene-play-v0 55 $\pm 1$ $42$ $\pm 4$ 51{51} $\pm 4$ 55 $\pm 1$ 1919 $\pm 2$ $38$ $\pm 3$ $\mathbf{64}$ $\pm 7$ $\mathbf{63}$ $\pm 6$
visual-antmaze visual-antmaze-medium-navigate-v0 $11$ $\pm 2$ $22$ $\pm 2$ $11$ $\pm 1$ 0 $\pm 0$ $\mathbf{94}$ $\pm 1$ $\mathbf{93}$ $\pm 4$ $55$ $\pm 47$ $\mathbf{95}$ $\pm 0$
visual-antmaze-large-navigate-v0 44 $\pm 0$ 55 $\pm 1$ 44 $\pm 1$ 0 $\pm 0$ $\mathbf{84}$ $\pm 1$ $53$ $\pm 9$ $43$ $\pm 44$ $\mathbf{82}$ $\pm 4$
visual-antmaze-giant-navigate-v0 0 $\pm 0$ 11 $\pm 1$ 0 $\pm 0$ 0 $\pm 0$ $\mathbf{47}$ $\pm 2$ 66 $\pm 4$ 44 $\pm 1$ $10$ $\pm 2$
visual-cube visual-cube-single-play-v0 55 $\pm 1$ $60$ $\pm 5$ $30$ $\pm 5$ $41$ $\pm 15$ $31$ $\pm 15$ $\mathbf{89}$ $\pm 0$ $63$ $\pm 37$ $\mathbf{88}$ $\pm 3$
visual-cube-double-play-v0 11 $\pm 1$ $10$ $\pm 2$ 11 $\pm 1$ 55 $\pm 0$ 22 $\pm 1$ $\mathbf{39}$ $\pm 2$ $28$ $\pm 6$ $\mathbf{40}$ $\pm 3$
visual-cube-triple-play-v0 $15$ $\pm 2$ $14$ $\pm 2$ $15$ $\pm 1$ $16$ $\pm 1$ $17$ $\pm 2$ $\mathbf{21}$ $\pm 0$ $18$ $\pm 1$ $\mathbf{20}$ $\pm 1$
visual-scene visual-scene-play-v0 $12$ $\pm 2$ $25$ $\pm 3$ $12$ $\pm 2$ $10$ $\pm 1$ $11$ $\pm 2$ $\mathbf{49}$ $\pm 4$ $38$ $\pm 3$ $\mathbf{47}$ $\pm 6$